### v2

# SemEval-2024 Task 6: SHROOM, a Shared-task on Hallucinations and Related Observable Overgeneration Mistakes

Abstract
This paper presents the results of the SHROOM, a shared task focused on detecting hallucinations: outputs from natural language generation (NLG) systems that are fluent, yet inaccurate. Such cases of overgeneration put in jeopardy many NLG applications, where correctness is often mission-critical. The shared task was conducted with a newly constructed dataset of 4000 model outputs labeled by 5 annotators each, spanning 3 NLP tasks: machine translation, paraphrase generation and definition modeling.
The shared task was tackled by a total of 58 different users grouped in 42 teams, out of which 27 elected to write a system description paper; collectively, they submitted over 300 prediction sets on both tracks of the shared task. We observe a number of key trends in how this approach was tackled—many participants rely on a handful of model, and often rely either on synthetic data for fine-tuning or zero-shot prompting strategies. While a majority of the teams did outperform our proposed baseline system, the performances of top-scoring systems are still consistent with a random handling of the more challenging items.
1 Introduction
The modern NLG landscape is plagued by two interlinked problems: On the one hand, our current neural models have a propensity to produce inaccurate but fluent outputs; on the other hand, our metrics are most apt at describing fluency, rather than correctness. This leads neural networks to “hallucinate”, i.e., produce fluent but incorrect outputs that we currently struggle to detect automatically. For instance, Dopierre et al. (2021) report that when trying to produce a paraphrase for the input “I am not sure where my phone is”, they obtain the following ‘hallucination’ behavior: “How can I find the location of any Android mobile”. For many NLG applications, the correctness of an output is however mission critical. For instance, producing a plausible-sounding translation that is inconsistent with the source text puts in jeopardy the usefulness of a machine translation pipeline.
Figure 1: The SHROOM logo.
This motivates us to organize a Shared-task on Hallucinations and Related Observable Overgeneration Mistakes, or SHROOM. With our shared task, we hope to foster the growing interest in this topic in the community (e.g., Ji et al., 2023; Raunak et al., 2021; Guerreiro et al., 2023; Xiao and Wang, 2021; Guo et al., 2022). In particular, in the SHROOM we adopt a post hoc setting, where models have already been trained and outputs already produced. Participants were asked to perform binary classification to identify cases of fluent overgeneration hallucinations in two different setups: model-aware and model-agnostic tracks. That is, participants had to detect grammatically sound outputs which contain incorrect or unsupported semantic information, inconsistent with the reference input, with or without having access to the model that produced the output.
To that end, we constructed a dataset comprising a collection of checkpoints, inputs, references and outputs of systems covering three different NLG tasks: definition modeling (DM, Noraset et al., 2017), machine translation (MT) and paraphrase generation (PG) trained with varying degrees of accuracy. Datapoints were all annotated by 5 human annotators each resulting in 1000 validation items and 3000 test items.
Beyond simply detecting factually unsupported outputs, one of the goals of this shared task was to establish whether hallucinations are best construed as a categorical phenomenon or a gradient one. Similar remarks have been made with respect to textual entailment (Bowman et al., 2015). As such, participants’ submission were scored both for accuracy (whether classifiers correctly identify hallucinations) and calibration (whether classifiers are confident about their prediction when they ought to be).
The shared task attracted a total of 58 different users grouped in 42 teams, out of which 27 elected to write a system description paper. Collectively, over the three weeks of the evaluation phase, participants submitted 300 valid sets of predictions on the model-aware track, and 320 on the model-agnostic track. We take this participation rate, along with the breadth of methodological approaches developed by participants, as clear signs of success for our shared task: This large pool of participants allows us to identify and discuss some key trends in how the task was tackled. Crucially, many participants rely on a handful of model, and often rely either on synthetic data for fine-tuning or zero-shot prompting strategies. In terms of raw performance, we note that while a majority of the teams ($64$ to $71\%$) did outperform our proposed baseline system, the performances of top-scoring systems are still consistent with a random handling of the more challenging items. In sum, this first iteration of the SHROOM underscores both an interest of the research community as well as the current limitations in our approaches.
The remainder of this article is structured as follows: Iin Section 2, we provide an overview of the current research landscape. Section 3 defines our theoretical framework, and Section 4 summarizes our data collection process. We then present and discuss shared task results in Sections 5 and 6 before concluding with a few thoughts on further research in Section 7.
2 Connecting with the past: related works and state of the art
It is now widely accepted that NLG models often generate outputs that are not faithful to the given input, commonly referred to in the community as hallucinations (Vinyals and Le, 2015; Raunak et al., 2021; Maynez et al., 2020). Yet there is minimal consensus on the optimal framework for its application. This lack of agreement is due in part to the diversity of tasks that NLG encompasses (Ji et al., 2023).
Guerreiro et al. (2023) propose a taxonomy of hallucinations that includes oscillatory productions, and fluent but strongly or fully “detached” outputs. While this taxonomy is well constructed, we find it inadequate for the needs of the community at large for four reasons: (i) It conflates some issues of fluency with semantic correctness (oscillatory productions are cases of non-fluent overgeneration where no extraneous semantic material is introduced); (ii) It only considers the most extreme cases of hallucinations (strongly or fully detached productions), whereas diagnosis of intermediary cases is bound to be more challenging and useful to the community; (iii) It focuses only on MT, although other tasks are also known to suffer from fluent overgeneration (e.g., Rohrbach et al., 2018), including the ones we propose to address; (iv) It uses only lowest scoring outputs, whereas any tool built to verify system outputs ought not to flag non-pathological outputs.
Alternative studies have built benchmarks for hallucination detection, with a predominant emphasis on dialogue systems. Li et al. (2023) propose the HaluEval benchmark using an annotation framework that does not necessarily center on the input given to the model and requires the annotators to search the internet for facts. Moreover, they opted to annotate the outputs of a popular LLM, with the major downsides that it is closed, not-transparent and commercial; rendering the research outputs that may stem from future studies less interesting. Other benchmarks include the works of Liu et al. (2022) and Zhou et al. (2021), which automatically insert hallucinations into training instances to generate syntactic data for token-level hallucination detection; Lin et al. (2022), which work with factual claims supported by reliable, publicly available evidence; and Dziri et al. (2022), which focus on knowledge-based dialogue systems and base their annotation on NLI, relying only on the system’s input, just as we do.
3 Tripping over hallucinations: task definition and annotation
In contrast with previous works (e.g. Guerreiro et al., 2023; Li et al., 2023), we focus on cases of fluent overgeneration since judgments pertaining to the over-generative nature of a production can be elicited by means of inferential semantics: if an output cannot be inferred from its semantic reference, then it contains some information that is not present in the reference—i.e., the model has generated more than we expected.111 Note that if the output can be inferred from the reference but the information is not explicitly present in the reference, then the model is actually making a correct semantic inference: it is generating a semantically sound output. E.g., if the the model produces “my tie is blue” for the reference “my tie is the color of the sky”, the model output is semantically sound. This approach connects with the theoretical framework sketched by van Deemter (2024), who likewise relies on inferential semantics but also considers undergeneration issues in NLG outputs. We provide multiple annotations and a gold majority label, given the low consensus on semantic annotations (Nie et al., 2020).
[…]
[ [ DM MT PG ].Model-agnostic [ DM MT PG ].Model-aware ] .SHROOM
Figure 2: Shared task overview. Both tracks feature all three NLG tasks. Datapoints from systems in blue correspond to target-referential datapoints and in red the ones that are either target- or source-referential; which we refer to as dual-referential.
In Figure 2 we provide an overview of the task. The SHROOM is framed around two key distinctions: (i) model-aware vs. model-agnostic approaches, and (ii) source-referential vs. dual-referential datapoints. The former corresponds to whether participants have access to the model that generated the item: Model-agnostic approaches are practical, as models may not be accessible to end users; Model-aware approaches can lead to richer and more accurate diagnoses. The latter is a consequence of our inferential take on over-generation: what can effectively serve as a semantic reference varies across NLP systems. For DM, where we fine-tune a language model to produce a definition for a given example of usage the datapoints are target-referential, i.e. the target is the sole usable semantic reference. In this context, the target serves as the sole usable semantic reference. Conversely, the target is expected to be semantically implied from the source in source-referential tasks, such as summarization. Note that we do not annotate source-referential tasks due to annotation challenges that make them unreliable for our purposes. In dual-referential tasks like PG & MT, this distinction bears no weight.
{ "hyp":"A cigarette .",
  "ref":"tgt",
  "src":"I stepped outside to smoke myself a j .
         What is the meaning of J ?",
  "tgt":"( plural Js or J s ) A marijuana
          cigarette .",
  "model":"ltg\/flan-t5-definition-en-base",
  "task":"DM",
  "labels":["Hallucination","Not Hallucination",
           "Not Hallucination", "Hallucination",
           "Hallucination"],
  "label":"Hallucination",
  "p(Hallucination)":0.6  }
Figure 3: Target-referential datapoint example from the validation set for the model-aware track.
In  Figure 3, we present an example datapoint displaying how we plan to encode all relevant information in a JSON format is provided. The datapoint keeps track of the source provided to the model as input (src), the intended target (tgt), the model production (hyp), the task this production was derived from (task), can correspond to DM, MT or PG), whether this datapoint is target-referential (ref), the annotations, the gold label and the proportion of annotators that labeled the utterance as a hallucination (labels, label, and p(Hallucination)). In the model-aware track, we will also provide a HuggingFace model name (model).
4 Foraging and harvesting season: Collected data
All SHROOM data (models, outputs and annotations) are available under a CC-BY license. [^2] Link to be added.
4.1 Data & model provenance
Participants have access to generated outputs from multiple systems trained to generate English output at various stages of their training, stemming from three sequence-to-sequence NLG tasks: DM, MT and PG. The SHROOM dataset consists of annotated test and dev sets, as well as a unlabeled training split of 30k datapoints per track and the full set of possible target references to allow corpus-wide approaches. To ensure effective annotation of the development and test sets, and to be able to guarantee a gradient in quality as measured by automated metrics, we pre-selected fluent outputs for the annotators, which we describe in the following. [^3] Note that we do not warranty that the training split contains fluent outputs.
MT:
For the model agnostic track we use the models from Mickus and Vázquez (2023). We compute perplexity for the all MT outputs and BERTScores with regards to the outputs and corresponding targets. We filter outputs with perplexity scores above the 2% quantile. From the filtered outputs, we randomly select 200 samples with BERTscores in the 1/7, 2/7, 3/7, 4/7, and 5/7 quantiles. For the model-aware track, we use the NLLB model (NLLB Team et al., 2022) and produce translations on the Flores-200 dataset from languages marked as low-resource to English. Next, we manually select a sample that is sufficiently fluent.
DM:
We use the model of Segonne and Mickus (2023) for the model-agnostic track, and for the model-aware track we used the flan-t5-definition-en-base (Giulianelli et al., 2023). We generate outputs on the English portion of the CoDWoE dataset (Mickus et al., 2022), and manually select a sample that is reasonably fluent and contains no profanities.
PG:
We used a pretrained and fine-tuned paraphrasing model [^4] https://huggingface.co/tuner007/pegasus_paraphrase based on Pegasus (Zhang et al., 2020) for the model-aware track, and the controlled paraphrase generation model of Vahtola et al. (2023) for the model-agnostic track.
We generated paraphrase hypotheses using Europarl (Koehn, 2005) and Opusparcus (Creutz, 2018) for the model-aware and -agnostic tracks, respectively. For the model-aware setup, we generated 50 hypotheses for each source sentence using diverse beam search (Vijayakumar et al., 2016) using BLEU scores (Papineni et al., 2002) to select the least similar hypothesis for each source sentence to serve as its paraphrase. For the model-agnostic setup, we calculated control tokens for each source sentence as in Vahtola et al. (2023), scaled the length-controlling value in range (1, 1.5) with a uniform probability distribution to provoke hallucination in the generated sequences, and used beam search with a beam size of 5 to produce the paraphrases. We manually curated the final validation and test examples.
4.2 Annotation
We annotate a total of 4,000 items, which are split 25%–75% between development and test sets: 1000 datapoints come from PG, 1500 from DM and 1500 from MT. Each item is annotated by five annotators on whether the reference entails the output. Annotations are binary, for ease of dataset construction. Gold labels are defined with respect to the annotators’ majority vote.
The annotators were enlisted via Prolific, [^5] https://www.prolific.com/ a paid platform specialized in gathering human data for research studies and AI dataset creation, among other purposes. We did not target any particular group of participants; the only screening prerequisites were that (i) participants had to be fluent in English and (ii) they should not have taken part in an initial pilot study.
We used Potato (Pei et al., 2022), an open-source annotation tool specifically designed to seamlessly integrate with Prolific. Annotators were first presented with a pre-annotation screen outlining the annotation guidelines, after which they commenced the annotation of items individually. Each item consisted of the Reference, the AI-generated output, and relevant context regarding the NLG task (DM, MT, or PG). The annotators were asked to answer the question "Does the following AI output only contain information supported by the Reference?" responding with either "yes" or "no," and were also given the opportunity to provide comments if necessary. Additionally, they could navigate back and forth through their assigned items. We set up a timer that notified the participants every 60 seconds of the time spent on an item. In Appendix B, we present a copy of the instructions we used.
To control for annotation quality, we manually reviewed annotations from two sets of selected annotators: (i) five randomly selected annotators; and (ii) the five annotators who completed the task the fastest (under 3.5 minutes). All 10 annotators completed 20 annotations each. We judged all 200 annotations to be sound, in that a reasoning could be reconstructed to explain the provided annotation.
(a) Validation split
(b) Test split
Figure 4: Distributions of annotations
Label distribution.
Figure 4 provides an overview of the distribution of labels in the SHROOM dataset splits (validation and test), broken down per NLG task (MT, DM and PG) and track (model-aware vs. model-agnostic). In this figure, we consider the empirical probability that a given item is judged to be a hallucination, i.e., the proportion of annotators judging the NLG output is not supported by the intended semantic reference.
[…]
Second, it is difficult to find hallucinations: The higher the empirical probability, the fewer the datapoints. This is especially true in the PG task: these outputs rarely yields consensual hallucinations, whereas we can find such items in DM and MT much more frequently. Looking at the expected value of the empirical probability per task, we find that DM consistently ranks higher than MT, which in turns ranks higher than PG. Both of these differences are significant under a one-sided Mann-Whitney U-test in the two test tracks ($p<0.0003$); in the model-aware validation dataset, only the difference between MT and PG is significant ($p<2\cdot 10^{-8}$), in the model-agnostic validation dataset, only the difference between DM and MT is ($p<0.04$). We note that DM requires a more complex processing of its input, as it has to rely on facts captured by the underlying LLM during its pre-training phase; for MT and PG, the input of the NLG task contains the semantic information necessary to produce a valid output. As such, we conjecture that the difficulty of an NLG task fosters hallucinatory behavior.666 We also remark that the two tracks are broadly comparable in terms of hallucinatory content. Two-samples Kolmogorov-Smirnov tests for either split (test or validation) do not provide sufficient grounds to suggest a difference of distribution in labels between model-aware and model-agnostic tracks—which again suggests that the relevant difference is at the task level, rather than at the model level.
5 They got so high: shared task results
The competition was held via Codalab (Pavao et al., 2023). The leaderboard was left hidden during the evaluation phase (i.e., participants were not notified of their submissions’ scores until the end of the evaluation phase) but users were allowed to make a high number of submissions (50).
Systems are evaluated according to two criteria: the accuracy that the system reached on the binary classification, and their calibration, measured as the Spearman correlation of the systems’ output probabilities with the proportion of the annotators marking the item as overgenerating. We rank systems by accuracy and break possible ties using calibration.
5.1 Baseline system
As a baseline for the task, we use an LLM [^7] We use quantized Mistral-7B-Instruct-v0.2 Jiang et al. (2023), from the Hugging Face hub huggingface.co/TheBloke/Mistral-7B-Instruct-v0.2-GGUF or the llamacpp project github.com/ggerganov/llama.cpp. to evaluate whether the generated hypotheses are coherent with the provided context. Drawing upon Manakul et al. (2023), we use the prompt template listed in Figure 5. The system of Manakul et al. (2023), which has gathered some attention from the community, constitutes a straightforward approach based on a modern LLM, and is therefore well-suited to serve as a baseline in our shared-task: it corresponds to a reasonable default approach to tackle the problem we challenge participants with.
Context: {}
Sentence: {}
Is the Sentence supported by the Context above?
Answer using ONLY yes or no:
Figure 5: Prompt template used in the baseline system, adapted from Manakul et al. (2023).
The specific context varies depending on the task addressed, i.e. the source sentence for the paraphrase generation task, and the target sentence for machine translation and definition modeling tasks. As for the probability of hallucination, we rely on the probability assigned by the model to the first output word. [^8] We note that this simple heuristic may not accurately represent the true hallucination probability. In cases where the output does not clearly indicate yes or no, we randomly select one, attributing a hallucination probability of $0.5$.
On the model-agnostic track, our baseline system achieves an accuracy of $0.697$ (with a calibration of $\rho=0.403$), on the model-aware track, we observe an accuracy of $0.745$ (with $\rho=0.488$). We can also indicate some other simple heuristics, such as picking the most frequent label (viz., Not Hallucination): In this case, one would expect an accuracy of $0.593$ on the model-agnostic track, and $0.633$ on the model-aware track. A purely random guess between the two possible labels would result in an accuracy of $0.5$. In short, our baseline systems systematically outperforms these crude heuristics.
5.2 Participating teams
A total of 59 individual users grouped in 42 teams participated in the shared task, out of which 27 elected to write a system description paper. During the evaluation phase, we received a total of 512 submissions, out of which 368 were successful. 264 of these submissions targeted both tracks, while 68 only targeted the model-agnostic track, and 36 only targeted the model-aware track. That is, we received 332 model-agnostic submissions and 300 model-aware submissions.
team Acc $\rho$
1 Halu-NLP (Mehta et al., 2024) 0.847 3330.8473330.847\,3330.847 333 0.769 5120.7695120.769\,5120.769 512
2 OPDAI (Chen et al., 2024) 0.836 $0$.8360000.836\,$0$.836 000 0.732 1950.7321950.732\,1950.732 195
3 HIT-MI&T Lab (Liu et al., 2024) 0.830 6670.8306670.830\,6670.830 667 0.767 7000.7677000.767\,7000.767 700
4 SHROOM-INDElab (Allen et al., 2024) 0.829 3330.8293330.829\,3330.829 333 0.721 0040.7210040.721\,0040.721 004
5 Alejandro Mosquera 0.826 $0$.8260000.826\,$0$.826 000 0.709 1720.7091720.709\,1720.709 172
6 DeepPavlov (Maksimov et al., 2024) 0.820 6670.8206670.820\,6670.820 667 0.752 4950.7524950.752\,4950.752 495
7 BruceW 0.820 6670.8206670.820\,6670.820 667 0.735 1790.7351790.735\,1790.735 179
8 TU Wien (Arzt et al., 2024) 0.816 6670.8166670.816\,6670.816 667 0.737 1380.7371380.737\,1380.737 138
9 SmurfCat (Rykov et al., 2024) 0.814 $0$.8140000.814\,$0$.814 000 0.723 4730.7234730.723\,4730.723 473
10 HaRMoNEE (Obiso et al., 2024) 0.814 $0$.8140000.814\,$0$.814 000 0.625 8800.6258800.625\,8800.625 880
11 AMEX AI LABS 0.812 6670.8126670.812\,6670.812 667 0.727 6820.7276820.727\,6820.727 682
12 Pollice Verso (Kobs et al., 2024) 0.802 6670.8026670.802\,6670.802 667 0.676 0240.6760240.676\,0240.676 024
13 MALTO (Borra et al., 2024) 0.800 6670.8006670.800\,6670.800 667 0.681 3080.6813080.681\,3080.681 308
14 UCC-NLP 0.795 3330.7953330.795\,3330.795 333 0.664 2440.6642440.664\,2440.664 244
15 Team CentreBack 0.792 $0$.7920000.792\,$0$.792 000 0.622 9220.6229220.622\,9220.622 922
16 Atresa 0.788 $0$.7880000.788\,$0$.788 000 0.645 6330.6456330.645\,6330.645 633
17 ustc_xsong 0.785 3330.7853330.785\,3330.785 333 0.694 8050.6948050.694\,8050.694 805
18 IRIT-Berger-Levrault (Bendahman et al., 2024) 0.782 6670.7826670.782\,6670.782 667 0.636 3240.6363240.636\,3240.636 324
19 silk_road 0.780 6670.7806670.780\,6670.780 667 0.671 5990.6715990.671\,5990.671 599
20 AILS NTUA (Griogoriadou et al., 2024) 0.778 $0$.7780000.778\,$0$.778 000 0.667 7600.6677600.667\,7600.667 760
21 zhuming 0.773 3330.7733330.773\,3330.773 333 0.480 9440.4809440.480\,9440.480 944
22 SibNN 0.770 $0$.7700000.770\,$0$.770 000 0.612 7590.6127590.612\,7590.612 759
23 UMUTeam (Pan et al., 2024) 0.769 3330.7693330.769\,3330.769 333 0.560 9460.5609460.560\,9460.560 946
24 NootNoot (Bahad et al., 2024) 0.764 6670.7646670.764\,6670.764 667 0.583 8350.5838350.583\,8350.583 835
25 HalluSafe (Rahimi et al., 2024) 0.763 3330.7633330.763\,3330.763 333 0.628 8700.6288700.628\,8700.628 870
26 Maha Bhaashya (Bhamidipati et al., 2024) 0.748 6670.7486670.748\,6670.748 667 0.605 2550.6052550.605\,2550.605 255
27 DUTh (Iordanidou et al., 2024) 0.744 $0$.7440000.744\,$0$.744 000 0.474 8320.4748320.474\,8320.474 832
28 Compos Mentis (Das and Srihari, 2024) 0.738 $0$.7380000.738\,$0$.738 000 0.595 0550.5950550.595\,0550.595 055
29 daixiang 0.737 3330.7373330.737\,3330.737 333 0.582 7410.5827410.582\,7410.582 741
30 NU-RU (Markchom et al., 2024) 0.728 $0$.7280000.728\,$0$.728 000 0.595 1260.5951260.595\,1260.595 126
baseline system 0.696 666 60.69666660.696\,666\,60.696 666 6 0.402 988 089 140 146 830.402988089140146830.402\,988\,089\,140\,146\,830.402 988 089 140 146 83
31 SLPL SHROOM (Fallah et al., 2024) 0.694 $0$.6940000.694\,$0$.694 000 0.423 2130.4232130.423\,2130.423 213
32 Skoltech 0.684 $0$.6840000.684\,$0$.684 000 0.674 3080.6743080.674\,3080.674 308
33 CAISA 0.676 6670.6766670.676\,6670.676 667 0.429 702-0.429702-0.429\,702- 0.429 702
34 AlphaIntellect (Choudhury et al., 2024) 0.654 $0$.6540000.654\,$0$.654 000 0.294 6080.2946080.294\,6080.294 608
35 deema 0.646 $0$.6460000.646\,$0$.646 000 0.565 6240.5656240.565\,6240.565 624
36 BrainLlama (Siino, 2024) 0.624 6670.6246670.624\,6670.624 667 0.203 6360.2036360.203\,6360.203 636
37 Byun (Byun, 2024) 0.616 6670.6166670.616\,6670.616 667 0.238 9790.2389790.238\,9790.238 979
38 Bolaca (Rösener et al., 2024) 0.612 6670.6126670.612\,6670.612 667 0.217 0620.2170620.217\,0620.217 062
most frequent guess 0.592 6670.5926670.592\,6670.592 667
39 AI Blues 0.587 3330.5873330.587\,3330.587 333 0.025 1040.0251040.025\,1040.025 104
random guess $0.5$
40 MARiA (Sanayei et al., 2024) 0.498 $0$.4980000.498\,$0$.498 000 0.025 0780.0250780.025\,0780.025 078
41 0x.Yuan (Lu, 2024) 0.460 6670.4606670.460\,6670.460 667 0.133 7200.1337200.133\,7200.133 720
(a) Model-agnostic track rankings
team Acc $\rho$ 1 HaRMoNEE (Obiso et al., 2024) 0.812 6670.8126670.812\,6670.812 667 0.699 3160.6993160.699\,3160.699 316 2 Halu-NLP (Mehta et al., 2024) 0.806 $0$.8060000.806\,$0$.806 000 0.714 7120.7147120.714\,7120.714 712 3 TU Wien (Arzt et al., 2024) 0.806 $0$.8060000.806\,$0$.806 000 0.707 1920.7071920.707\,1920.707 192 4 OPDAI (Chen et al., 2024) 0.805 3330.8053330.805\,3330.805 333 0.680 2730.6802730.680\,2730.680 273 5 HIT-MI&T Lab (Liu et al., 2024) 0.804 6670.8046670.804\,6670.804 667 0.712 3250.7123250.712\,3250.712 325 6 SHROOM-INDElab (Allen et al., 2024) 0.802 $0$.8020000.802\,$0$.802 000 0.655 6700.6556700.655\,6700.655 670 7 AMEX AI LABS 0.800 6670.8006670.800\,6670.800 667 0.695 7100.6957100.695\,7100.695 710 8 DeepPavlov (Maksimov et al., 2024) 0.799 3330.7993330.799\,3330.799 333 0.712 6930.7126930.712\,6930.712 693 9 silk_road 0.798 $0$.7980000.798\,$0$.798 000 0.686 5630.6865630.686\,5630.686 563 10 AILS NTUA (Griogoriadou et al., 2024) 0.794 6670.7946670.794\,6670.794 667 0.685 1960.6851960.685\,1960.685 196 11 BruceW 0.794 $0$.7940000.794\,$0$.794 000 0.659 9600.6599600.659\,9600.659 960 12 Team CentreBack 0.789 3330.7893330.789\,3330.789 333 0.606 3040.6063040.606\,3040.606 304 13 UCC-NLP 0.788 6670.7886670.788\,6670.788 667 0.644 0660.6440660.644\,0660.644 066 14 ustc_xsong 0.787 3330.7873330.787\,3330.787 333 0.657 9690.6579690.657\,9690.657 969 15 UMUTeam (Pan et al., 2024) 0.784 $0$.7840000.784\,$0$.784 000 0.506 8950.5068950.506\,8950.506 895 16 HalluSafe (Rahimi et al., 2024) 0.783 3330.7833330.783\,3330.783 333 0.537 4020.5374020.537\,4020.537 402 17 SmurfCat (Rykov et al., 2024) 0.782 6670.7826670.782\,6670.782 667 0.671 2000.6712000.671\,2000.671 200 18 Atresa 0.782 6670.7826670.782\,6670.782 667 0.623 9140.6239140.623\,9140.623 914 19 IRIT-Berger-Levrault (Bendahman et al., 2024) 0.781 3330.7813330.781\,3330.781 333 0.601 2760.6012760.601\,2760.601 276 20 Pollice Verso (Kobs et al., 2024) 0.777 3330.7773330.777\,3330.777 333 0.600 7360.6007360.600\,7360.600 736 21 NU-RU (Markchom et al., 2024) 0.768 $0$.7680000.768\,$0$.768 000 0.581 9070.5819070.581\,9070.581 907 22 zhuming 0.768 $0$.7680000.768\,$0$.768 000 0.471 5990.4715990.471\,5990.471 599 23 SibNN 0.762 6670.7626670.762\,6670.762 667 0.586 9430.5869430.586\,9430.586 943 24 Compos Mentis (Das and Srihari, 2024) 0.756 $0$.7560000.756\,$0$.756 000 0.565 6980.5656980.565\,6980.565 698 25 DUTh (Iordanidou et al., 2024) 0.755 3330.7553330.755\,3330.755 333 0.528 4800.5284800.528\,4800.528 480 baseline system 0.745 333 333 333 333 30.74533333333333330.745\,333\,333\,333\,333\,30.745 333 333 333 333 3 0.487 886 768 695 713 960.487886768695713960.487\,886\,768\,695\,713\,960.487 886 768 695 713 96 26 AlphaIntellect (Choudhury et al., 2024) 0.711 3330.7113330.711\,3330.711 333 0.426 4290.4264290.426\,4290.426 429 27 SLPL SHROOM (Fallah et al., 2024) 0.706 $0$.7060000.706\,$0$.706 000 0.426 2270.4262270.426\,2270.426 227 28 deema 0.688 $0$.6880000.688\,$0$.688 000 0.519 1040.5191040.519\,1040.519 104 29 BrainLlama (Siino, 2024) 0.671 3330.6713330.671\,3330.671 333 0.243 7660.2437660.243\,7660.243 766 30 daixiang 0.649 3330.6493330.649\,3330.649 333 0.217 6700.2176700.217\,6700.217 670 most frequent guess 0.632 6670.6326670.632\,6670.632 667 31 Bolaca (Rösener et al., 2024) 0.626 $0$.6260000.626\,$0$.626 000 0.283 2360.2832360.283\,2360.283 236 32 NootNoot (Bahad et al., 2024) 0.612 6670.6126670.612\,6670.612 667 0.355 0610.3550610.355\,0610.355 061 33 Byun (Byun, 2024) 0.610 $0$.6100000.610\,$0$.610 000 0.234 0480.2340480.234\,0480.234 048 34 Maha Bhaashya (Bhamidipati et al., 2024) 0.606 $0$.6060000.606\,$0$.606 000 0.209 2350.2092350.209\,2350.209 235 35 CAISA 0.567 3330.5673330.567\,3330.567 333 0.100 493-0.100493-0.100\,493- 0.100 493 36 Skoltech 0.557 3330.5573330.557\,3330.557 333 0.011 140-0.011140-0.011\,140- 0.011 140 37 MARiA (Sanayei et al., 2024) 0.505 3330.5053330.505\,3330.505 333 0.008 9850.0089850.008\,9850.008 985 random guess $0.5$ 38 octavianB (Brodoceanu, 2024) 0.483 3330.4833330.483\,3330.483 333 0.063 792-0.063792-0.063\,792- 0.063 792 team Acc $\rho$ 1 Halu-NLP (Mehta et al., 2024) 0.847 3330.8473330.847\,3330.847 333 0.769 5120.7695120.769\,5120.769 512 2 OPDAI (Chen et al., 2024) 0.836 $0$.8360000.836\,$0$.836 000 0.732 1950.7321950.732\,1950.732 195 3 HIT-MI&T Lab (Liu et al., 2024) 0.830 6670.8306670.830\,6670.830 667 0.767 7000.7677000.767\,7000.767 700 4 SHROOM-INDElab (Allen et al., 2024) 0.829 3330.8293330.829\,3330.829 333 0.721 0040.7210040.721\,0040.721 004 5 Alejandro Mosquera 0.826 $0$.8260000.826\,$0$.826 000 0.709 1720.7091720.709\,1720.709 172 6 DeepPavlov (Maksimov et al., 2024) 0.820 6670.8206670.820\,6670.820 667 0.752 4950.7524950.752\,4950.752 495 7 BruceW 0.820 6670.8206670.820\,6670.820 667 0.735 1790.7351790.735\,1790.735 179 8 TU Wien (Arzt et al., 2024) 0.816 6670.8166670.816\,6670.816 667 0.737 1380.7371380.737\,1380.737 138 9 SmurfCat (Rykov et al., 2024) 0.814 $0$.8140000.814\,$0$.814 000 0.723 4730.7234730.723\,4730.723 473 10 HaRMoNEE (Obiso et al., 2024) 0.814 $0$.8140000.814\,$0$.814 000 0.625 8800.6258800.625\,8800.625 880 11 AMEX AI LABS 0.812 6670.8126670.812\,6670.812 667 0.727 6820.7276820.727\,6820.727 682 12 Pollice Verso (Kobs et al., 2024) 0.802 6670.8026670.802\,6670.802 667 0.676 0240.6760240.676\,0240.676 024 13 MALTO (Borra et al., 2024) 0.800 6670.8006670.800\,6670.800 667 0.681 3080.6813080.681\,3080.681 308 14 UCC-NLP 0.795 3330.7953330.795\,3330.795 333 0.664 2440.6642440.664\,2440.664 244 15 Team CentreBack 0.792 $0$.7920000.792\,$0$.792 000 0.622 9220.6229220.622\,9220.622 922 16 Atresa 0.788 $0$.7880000.788\,$0$.788 000 0.645 6330.6456330.645\,6330.645 633 17 ustc_xsong 0.785 3330.7853330.785\,3330.785 333 0.694 8050.6948050.694\,8050.694 805 18 IRIT-Berger-Levrault (Bendahman et al., 2024) 0.782 6670.7826670.782\,6670.782 667 0.636 3240.6363240.636\,3240.636 324 19 silk_road 0.780 6670.7806670.780\,6670.780 667 0.671 5990.6715990.671\,5990.671 599 20 AILS NTUA (Griogoriadou et al., 2024) 0.778 $0$.7780000.778\,$0$.778 000 0.667 7600.6677600.667\,7600.667 760 21 zhuming 0.773 3330.7733330.773\,3330.773 333 0.480 9440.4809440.480\,9440.480 944 22 SibNN 0.770 $0$.7700000.770\,$0$.770 000 0.612 7590.6127590.612\,7590.612 759 23 UMUTeam (Pan et al., 2024) 0.769 3330.7693330.769\,3330.769 333 0.560 9460.5609460.560\,9460.560 946 24 NootNoot (Bahad et al., 2024) 0.764 6670.7646670.764\,6670.764 667 0.583 8350.5838350.583\,8350.583 835 25 HalluSafe (Rahimi et al., 2024) 0.763 3330.7633330.763\,3330.763 333 0.628 8700.6288700.628\,8700.628 870 26 Maha Bhaashya (Bhamidipati et al., 2024) 0.748 6670.7486670.748\,6670.748 667 0.605 2550.6052550.605\,2550.605 255 27 DUTh (Iordanidou et al., 2024) 0.744 $0$.7440000.744\,$0$.744 000 0.474 8320.4748320.474\,8320.474 832 28 Compos Mentis (Das and Srihari, 2024) 0.738 $0$.7380000.738\,$0$.738 000 0.595 0550.5950550.595\,0550.595 055 29 daixiang 0.737 3330.7373330.737\,3330.737 333 0.582 7410.5827410.582\,7410.582 741 30 NU-RU (Markchom et al., 2024) 0.728 $0$.7280000.728\,$0$.728 000 0.595 1260.5951260.595\,1260.595 126 baseline system 0.696 666 60.69666660.696\,666\,60.696 666 6 0.402 988 089 140 146 830.402988089140146830.402\,988\,089\,140\,146\,830.402 988 089 140 146 83 31 SLPL SHROOM (Fallah et al., 2024) 0.694 $0$.6940000.694\,$0$.694 000 0.423 2130.4232130.423\,2130.423 213 32 Skoltech 0.684 $0$.6840000.684\,$0$.684 000 0.674 3080.6743080.674\,3080.674 308 33 CAISA 0.676 6670.6766670.676\,6670.676 667 0.429 702-0.429702-0.429\,702- 0.429 702 34 AlphaIntellect (Choudhury et al., 2024) 0.654 $0$.6540000.654\,$0$.654 000 0.294 6080.2946080.294\,6080.294 608 35 deema 0.646 $0$.6460000.646\,$0$.646 000 0.565 6240.5656240.565\,6240.565 624 36 BrainLlama (Siino, 2024) 0.624 6670.6246670.624\,6670.624 667 0.203 6360.2036360.203\,6360.203 636 37 Byun (Byun, 2024) 0.616 6670.6166670.616\,6670.616 667 0.238 9790.2389790.238\,9790.238 979 38 Bolaca (Rösener et al., 2024) 0.612 6670.6126670.612\,6670.612 667 0.217 0620.2170620.217\,0620.217 062 most frequent guess 0.592 6670.5926670.592\,6670.592 667 39 AI Blues 0.587 3330.5873330.587\,3330.587 333 0.025 1040.0251040.025\,1040.025 104 random guess $0.5$ 40 MARiA (Sanayei et al., 2024) 0.498 $0$.4980000.498\,$0$.498 000 0.025 0780.0250780.025\,0780.025 078 41 0x.Yuan (Lu, 2024) 0.460 6670.4606670.460\,6670.460 667 0.133 7200.1337200.133\,7200.133 720
(b) Model-aware track rankings
Table 1: SHROOM team rankings. Codalab usernames are used to define teams when no other information was provided.
We present the model-agnostic track rankings in Table 1a and the model-aware track in Table 1b. As one might expect, there is a high correlation between the accuracy and calibration scores of each team’s top ranking submission, which translates into a Spearman’s $\rho$ correlation of $0.909$ on the model-agnostic track and $0.949$ on the model-aware track. Most of the top submissions per team rank above our baseline ($30/42\approx 71.4\%$ in the model-agnostic track, $25/39\approx 64.1\%$ in the model-aware track). This appears roughly in line with all submissions globally: $69.9\%$ of all model-agnostic submissions and $57.0\%$ of all model-aware submissions score higher than our baseline.
Another point worth stressing is that teams that fare well on one track usually fare equally well on the other: For the 38 teams participating in both tracks, we find that the rank they obtain on the model-aware track correlates with the rank they obtain on the model-agnostic track (Spearman’s $\rho=0.884$). This would tentatively suggest that participants could not effectively leverage the supplementary data available in the model-aware track. [^9] An alternative account would be that all teams that participated in both tracks equally benefited from the access to the model weights, which we deem much less likely.
Lastly, we note that there is a ceiling in terms of performances: The most effective systems misclassify between 15 to 19% of all items, or almost one in every six or five datapoints. We have discussed above that, as hallucinations are a graded phenomenon, a large segment of our data (30%) corresponds to ambiguous cases where annotators are split 2 vs. 3. As such, it is worth stressing that top scores are consistent with models that classify consensual items well (where at most one annotator disagree), but perform at random chance on the more challenging ambiguous datapoints.
6 A bunch of fun guys: qualitative analysis of participants systems
We derive our analyses from system description papers as well as self-reports from a handful of participants who elected to not provide a full description of their systems. This corresponds to 33 systems out of the 42 identified teams that participated to the shared task, out of which 6 did not provide a full description. See also Table 2 in Appendix A for further details.
Figure 6: Known models used by more than one team. A full circle would correspond to a given model used by all of respondents, half a circle to 50% of respondents using said model. Best viewed in color.
How the task was approached.
The teams used a variety of methods to address the problem, ranging from ensemble techniques to fine-tuning pretrained language models (LLMs) and prompt engineering. As expected, most teams used popular pre-trained LLMs such as GPT, LLaMA, DeBERTa, RoBERTa, and XLM-RoBERTa; Figure 6 provides a summary of which models were most popular among our teams. The Vectara hallucination evaluation model [^10] https://huggingface.co/vectara/hallucination_evaluation_model turned out to be extremely popular, as more than 1 in 4 teams that provided information about their systems report having used it in their experiments. If we add other DeBERTa-based models, this number climbs to $16/33$, i.e. almost every other team used DeBERTa or a variant thereof.
Yet, the ways in which these LLMs were used cover a wide range of approaches: Some either fine-tuned on hallucination data or optimized with prompts; others employed in-context learning with role-playing, automatic prompt generation, and ensemble methods. Furthermore, some teams focused on zero-shot and few-shot approaches, while others focused on synthetic data generation and semi-supervised learning techniques to construct a labeled training set. Especially noteworthy, Rahimi et al. (2024) report constructing a manual dataset of 3000 datapoints for training their systems.
Teams predominantly relied on the data constructed for the SHROOM, although some teams added datasets such as QQP and PAWS. Interestingly, we also note five teams relying on NLI/entailment data or models, including some that achieved high results (Obiso et al., 2024; Sanayei et al., 2024; Borra et al., 2024; Liu et al., 2024 and Team CentreBack)—and this matches the theoretical framework adopted in this shared task.
What worked well.
We now turn to what distinguishes top scorers from other submissions. We note that systems based on the closed-source models GPT-3.5 and GPT-4 tend to fare well: 4 out of the 6 highest scoring systems on either track—Mehta et al. (2024); Obiso et al. (2024); Liu et al. (2024); Allen et al. (2024) and Alejandro Mosquera—all report using these models. This is however not a strict requisite as OPDAI (Chen et al., 2024) manages to rank high (2nd on the model-agnostic track and 4th on the model-aware track) without it. Neither does using closed-source models guarantee a high result: UCC-NLP and Markchom et al. (2024) also use GPT-3.5, and while the former is ranked 14th on the model-agnostic track and 13th on the model-aware track, the latter is ranked 30th on the model-agnostic track and 21st on the model-aware track, and only outperforms the baseline model in accuracy by 0.02 to 0.03 points.
Remarkably, many of the top-scoring approaches rely on fine-tuning (Liu et al., 2024; Obiso et al., 2024; Arzt et al., 2024; Chen et al., 2024) or ensembling (Mehta et al., 2024; Maksimov et al., 2024, Alejandro Mosquera), suggesting that high performances do not come out of the box from off-the-shelf LLMs and systems. It is necessary to adapt existing models or properly establish to what extent their predictions are useful to the task at hand.
Figure 7: Rank obtained vs. number of submissions made on both tracks.
Another important trend we identify is that the number of submissions per team anti-correlates with the rank they obtain: The more participants submitted, the higher their best scores went. This is visualized in Figure 7: On both tracks, we find reasonable anti-correlations ($-0.58<\rho<-0.44$) indicating that top-scorers tended to submit more. This might provide an alternative explanation for what distinguishes top-scorers from other participants: If we were to model participants’ submissions as a random process, we would expect that sampling more often (i.e., submitting more) would mechanically yield a better rank.
Overall, the high methodological diversity highlights the complexity of hallucination detection, even when contained the simple inferential semantics framework of our shared task: While a focus on NLI or using high-performance closed source models may help, the highest scores are obtained through thorough involvement—both in terms of model training and prediction set submissions.
7 Much room to grow: conclusions and future perspectives
This first iteration of the SHROOM shared task on detecting hallucinations has allowed us to make significant headway into understanding the confabulatory behavior of modern NLG systems. The data collected demonstrate that hallucinations correspond to a gradient phenomenon, and that different speakers form different opinions as to what counts as a hallucination. We were also able to showcase that ambiguous items remain challenging, and that the current state of the art on the dataset we provided is compatible with simple random guesses whenever the data is more ambiguous. This results underscore the massive gap that NLP research urgently needs to address: one out of every six items is still misclassified by the most effective systems showcased during this shared task.
The diversity of methodologies employed by participants underscores how out-of-the-box solutions are not sufficient: Highest scoring teams had to rely on fine-tuning or ensembling and made a high number of submissions. Relatedly, access to the model parameters was of limited help: Few approaches attempted to perform model-specific investigations, and performances on the model-aware track are in fact lower than what we observed on the model-agnostic track. Properly leveraging the parameter space for finer-grained hallucination detection remains a point for future research to investigate.
This shared task has not broached some crucial aspects and questions: How do these results translate insofar as modern LLMs—often much larger and better trained than the systems we studied here—are concerned? Can we leverage sentence-level predictions to pinpoint token-level issues with the output of our NLG systems? And will the difficulties that we underscored in this purely English be exacerbated when studying other languages—especially those that are less well-resourced and typologically different? Answering these questions and more will require further research—and perhaps future iterations of this shared task.
Overall, the success of this shared task is owed to its committed participants. We received over 350 submissions in the span of three weeks from across the world. The width of approaches studied and reported upon provides a useful snapshot of where the field is at, what approaches are favored, and what gaps still need to be overcome. We expect that the results of the SHROOM will provide a useful starting point for future work on hallucinations.
Doing SHROOM responsibly: ethical considerations
We strive to adhere to the ACL Code of Ethics.
Broader Impact.
Hallucinated outputs from large language models can be used to further spread disinformation and advance misleading narratives. Detecting hallucinated outputs is an important step in elucidating the factors of this phenomena and contribute to ongoing efforts to mitigate hallucination. This leads to the development of more trustworthy generative language models.
Data and Annotators.
Our annotators were suitably compensated for their work in excess of minimum wage. Due to the nature of the proposed task, the data we release might contain false or misleading statements. In the case of annotated data, these statements are labeled as such, but this does not for the unannotated portions of the data. We manually pre-filtered the data to remove profanities before providing them to annotators. Such precautions were not taken for the unannotated portion of the dataset, which might therefore contain offensive, obscene or otherwise unconscionable items.
Acknowledgments
The construction of the SHROOM dataset was made possible by a grant from the Oskar Öfflund Foundation. This work is also supported by the ICT 2023 project “Uncertainty-aware neural language models” funded by the Academy of Finland (grant agreement № 345999). We also thank the CSC-IT Center for Science Ltd., for computational resources.
The shared task logo (cf. Figure 1) uses the “Retro Cool” font from Nirmana Visual (https://nirmanavisual.com/), made available for personal / non-commercial uses.
Appendix A Shared consciousnesses: Overview of approaches used by SHROOM teams
Team & Paper Resources Overview
AI Blues
(No report)
AILS NTUA
Griogoriadou et al. (2024)
SHROOM datasets; Vectara model.
Fine-tuned models and voting classifier.
Alejandro Mosquera
SHROOM datasets; COMET, Vectara, LaBSE, GPT35 and GPT4 models.
Ensemble of publicly available models. Logistic Regression was used as final scoring model.
AlphaIntellect
Choudhury et al. (2024)
SHROOM dataset, SBERT
Fully-connected neural network classifiers with SBERT embeddings as input.
AMEX AI LABS
SHROOM datasets; Vectara and OpenChat models.
Ensemble of LLM (using Openchat) zero shot and few shot with Vectara cross encoder based scores.
Atresa
(No report)
BrainLlama
Siino (2024)
LLaMA model.
Prompt-based approach with LLaMA.
BruceW
(No report)
Byun
Byun (2024)
SHROOM dataset, data augmentation, RoBERTa
Finetuned a BERT or RoBERTa model with a softmax layer to output the probability of hallucinated text. Finetuning data is the labelled SHROOM data augmented with data points constructed by replacing words with synonyms.
CAISA
(No report)
Compos Mentis
Das and Srihari (2024)
HalluEval dataset; Mistral 7B instruct model.
Ensemble of several role-based LLMs, which were either fine-tuned on hallucination data or role-based prompting.
daixiang
(No report)
deema
(No report)
DeepPavlov
Maksimov et al. (2024)
SHROOM dataset; OpenChat, DeBERTa, RoBERTa and T5 models.
Ensemble of several pretrained Transformer-based models to get features for validation and test data of SHROOM dataset and trained a boosting-based meta-model on top.
DUTh
Iordanidou et al. (2024)
SHROOM, LaBSE, T5, DistilUSE
Using pre-trained LLMs and classifiers
Bolaca
Rösener et al. (2024)
SHROOM dataset, SBERT
Logistic regression and feed-forward classifier trained on SBERT embeddings
HalluSafe
Rahimi et al. (2024)
SHROOM, labeled 3000 samples of the training data
Fine-tuned a DeBERTa-v3-large
Halu-NLP
Mehta et al. (2024)
SHROOM datasets; GPT, SelfCheckGPT and Vectara models.
Prompts and GroupCheckGPT. NB: due to a team name change, this team is also referred to as GroupCheckGPT by some participants.
HaRMoNEE
Obiso et al. (2024)
SHROOM, SNLI, MNLI and PAWS datasets; Vectara and GPT4 models.
Highest results obtained with zero-shot prompting in the model-aware track; pretraining on NLI and PAWS followed by finetuning on the model-agnostic track.
HIT-MI&T Lab
Liu et al. (2024)
SHROOM with training dataset labeled using GPT-4; DeBERTaV3, InternLM2, SBERT, and UniEval.
Fine-tune the DeBERTaV3 and InternLM2 models, and call the SBERT and UniEval models to select the optimal threshold usinf SHROOM & syntheticaly labeled data. The system obtians the final results by combining the prediction results of each model.
IRIT-Berger-Levrault
Bendahman et al. (2024)
SHROOM datasets; Sentence-t5, BGE, e5 models.
Computes the cosine similarity of sentence embeddings and classify based on an empirical threshold value.
Maha Bhaashya
Bhamidipati et al. (2024)
DeBERTa models.
Zero shot inference, pretrained cross encoder model
MALTO
Borra et al. (2024)
SHROOM model-agnostic dataset, DeBERTa pretrained and finetuned on MNLI, SOLAR-10.7B quantized from TheBloke (for synthetic data generation)
Encoder and classifier, fine-tuned in various ways (including with synthetic data)
MARiA
Sanayei et al. (2024)
SHROOM dataset, SBERT, bart-large-mnli, Mixtral
Three approaches: (1) Cosine similarity of SBERT embeddings between source-hypothesis and source-target pairs; (2) NLI classification using bart-large-mnli model; and (3) Mixtral prompting. Only the Mixtral results were submitted.
NootNoot
Bahad et al. (2024)
SHROOM dataset; Mixtral and RoBERTa models.
Mixtral prompting and RoBERTa finetuning.
NU-RU
Markchom et al. (2024)
SHROOM, GPT-3.5, Sentence Transformers
Tried two approaches: (1) hypothesis-target cosine similarity, using a threshold value to determine whether the hypothesis is a hallucation. (2) SelfCheckGPT with a customized prompt for each NLG task, designed to assess its coherence with the provided source and target. Each prompt is iterated through the GPT-3.5 model five times, and the final label is determined by the majority response.
octavianB
Brodoceanu (2024)
RoBERTa
Used a pretrained model (roberta-large-openai-detector) that has been trained to distinguish between text generated by LLMs and text written by humans.
OPDAI
Chen et al. (2024)
SHROOM, Mistral-7B-Instruct-v0.2, self constructed training data
Supervised fine-tuning over synthetically constructed weakly supervised training data.
Pollice Verso
Kobs et al. (2024)
Mistral2, LLaMa2, Phi2 and Zephyr models; uses SHROOM train set for prompt optimization.
Ensembling over the output logits of prompt-based LLMs (mistral, llama etc) after automatically optimizing their prompts ("OPRO").
SHROOM-INDElab
Allen et al. (2024)
SHROOM dataset; GPT 3.5 and GPT 4 models.
In-context learning with role-play and automatic prompt generation in a few-shot classfier, using a closed-source LLM.
SibNN
SHROOM datasets; XLM-RoBERTa model.
Fine-tunes a self-adaptive hierarchical variant of XLM-RoBERTa-XL twice: first as an embedder (in a few-shot mode), then as a binary classifier. More details at
https://huggingface.co/bond005/xlm-roberta-xl-hallucination-detector.
silk_road
SHROOM datasets; Vectara model.
Fine-tunes an off-the-shelf Cross-Encoder hallucination evaluation model.
Skoltech
(No report)
SLPL SHROOM
Fallah et al. (2024)
SHROOM datasets; LaBSE, DeBERTa, Zephyr, Mistral and Llama2 models.
Using two LLMs to classify and explain their decision and another LLM to judge and decide based on those explanations.
SmurfCat
Rykov et al. (2024)
SHROOM (synthetically augmented), QQP and PAWS datasets; E5, T5, Vectara models.
Fine-tuning of e5-mistral-7b-instruct using synthetic data collected with LLaMA2-7B adapters trained to produce data with and without hallucinations. However, there are two other systems: one works as a voting ensemble of multiple LLMs, and another uses the Mutual Implication Score architecture.
Team CentreBack
SHROOM dataset; DeBERTa model.
Uses an off-the-shelf library (SelfCheckGPT’s SelfCheckNLI function) to calculate contradiction scores on a small labeled test set and then defined a threshold for hallucination.
TU Wien
Arzt et al. (2024)
SHROOM dataset; Vectara model.
Model-aware track best submissions uses a Vectara hallucination detection model finetuned on the validation set. The best model-agnostic track submission is a meta-model that utilizes linear regression and is trained on features that correspond to probabilities predicted by individual systems we implemented.
UCC-NLP
SHROOM dataset; GPT-3.5 and Vectara models.
Uses BertScore and GPT-3.5 to create synthetic labels and fine-tune a Vectara LLM.
UMUTeam
Pan et al. (2024)
SHROOM dataset; TULU-DPO model.
Zero-shot approach
ustc_xsong
(No report)
zhuming
(No report)
0x.Yuan
Lu (2024)
Mistral, Mixtral, LLaMA, Falcon, WizardLM and Capybara models.
Zero-shot prompt engineering. Expects most LLMs will have different hallucination patterns, and tests whether ensembling can mitigate this.
Table 2: Participating teams and their respective works.
In Table 2, we provide a short overview of the various teams, the resources they utilized (models & datasets), as well as a short description of their approach.
Appendix B What SHROOM makes you do: Annotation guidelines
In Figure 8, we provide an exact copy of the annotation guidelines given to the annotators. These guidelines are based on five of the organizers’ experience of annotating the trial set, and were provided to annotators recruited for the validation and test splits.
Figure 8: Annotation guidelines.

### v3

@@
 4 Foraging and harvesting season: Collected data
-All SHROOM data (models, outputs and annotations) are available under a CC-BY license. [^2] Link to be added.
+All SHROOM data (models, outputs and annotations) are available under a CC-BY license. [^2] See helsinki-nlp.github.io/shroom
 4.1 Data & model provenance
@@
 Team & Paper Resources Overview
-AI Blues
-(No report)
+AI Blues (No report)
 AILS NTUA
 Griogoriadou et al. (2024)
-SHROOM datasets; Vectara model.
-Fine-tuned models and voting classifier.
-Alejandro Mosquera
-SHROOM datasets; COMET, Vectara, LaBSE, GPT35 and GPT4 models.
-Ensemble of publicly available models. Logistic Regression was used as final scoring model.
+SHROOM datasets; Vectara model. Fine-tuned models and voting classifier.
+Alejandro Mosquera SHROOM datasets; COMET, Vectara, LaBSE, GPT35 and GPT4 models. Ensemble of publicly available models. Logistic Regression was used as final scoring model.
 AlphaIntellect
 Choudhury et al. (2024)
-SHROOM dataset, SBERT
-Fully-connected neural network classifiers with SBERT embeddings as input.
-AMEX AI LABS
-SHROOM datasets; Vectara and OpenChat models.
-Ensemble of LLM (using Openchat) zero shot and few shot with Vectara cross encoder based scores.
-Atresa
-(No report)
+SHROOM dataset, SBERT Fully-connected neural network classifiers with SBERT embeddings as input.
+AMEX AI LABS SHROOM datasets; Vectara and OpenChat models. Ensemble of LLM (using Openchat) zero shot and few shot with Vectara cross encoder based scores.
+Atresa (No report)
 BrainLlama
 Siino (2024)
-LLaMA model.
-Prompt-based approach with LLaMA.
-BruceW
-(No report)
+LLaMA model. Prompt-based approach with LLaMA.
+BruceW (No report)
 Byun
 Byun (2024)
-SHROOM dataset, data augmentation, RoBERTa
-Finetuned a BERT or RoBERTa model with a softmax layer to output the probability of hallucinated text.
+SHROOM dataset, data augmentation, RoBERTa Finetuned a BERT or RoBERTa model with a softmax layer to output the probability of hallucinated text.
 Finetuning data is the labelled SHROOM data augmented with data points constructed by replacing words with synonyms.
-CAISA
-(No report)
+CAISA (No report)
 Compos Mentis
 Das and Srihari (2024)
-HalluEval dataset; Mistral 7B instruct model.
-Ensemble of several role-based LLMs, which were either fine-tuned on hallucination data or role-based prompting.
-daixiang
-(No report)
-deema
-(No report)
+HalluEval dataset; Mistral 7B instruct model. Ensemble of several role-based LLMs, which were either fine-tuned on hallucination data or role-based prompting.
+daixiang (No report)
+deema (No report)
 DeepPavlov
@@
 Iordanidou et al. (2024)
-SHROOM, LaBSE, T5, DistilUSE
-Using pre-trained LLMs and classifiers
+SHROOM, LaBSE, T5, DistilUSE Using pre-trained LLMs and classifiers
 Bolaca
 Rösener et al. (2024)
-SHROOM dataset, SBERT
-Logistic regression and feed-forward classifier trained on SBERT embeddings
+SHROOM dataset, SBERT Logistic regression and feed-forward classifier trained on SBERT embeddings
 HalluSafe
 Rahimi et al. (2024)
-SHROOM, labeled 3000 samples of the training data
-Fine-tuned a DeBERTa-v3-large
+SHROOM, labeled 3000 samples of the training data Fine-tuned a DeBERTa-v3-large
 Halu-NLP
 Mehta et al. (2024)
-SHROOM datasets; GPT, SelfCheckGPT and Vectara models.
-Prompts and GroupCheckGPT. NB: due to a team name change, this team is also referred to as GroupCheckGPT by some participants.
+SHROOM datasets; GPT, SelfCheckGPT and Vectara models. Prompts and GroupCheckGPT. NB: due to a team name change, this team is also referred to as GroupCheckGPT by some participants.
 HaRMoNEE
@@
 Bendahman et al. (2024)
-SHROOM datasets; Sentence-t5, BGE, e5 models.
-Computes the cosine similarity of sentence embeddings and classify based on an empirical threshold value.
+SHROOM datasets; Sentence-t5, BGE, e5 models. Computes the cosine similarity of sentence embeddings and classify based on an empirical threshold value.
 Maha Bhaashya
 Bhamidipati et al. (2024)
-DeBERTa models.
-Zero shot inference, pretrained cross encoder model
+DeBERTa models. Zero shot inference, pretrained cross encoder model
 MALTO
 Borra et al. (2024)
-SHROOM model-agnostic dataset, DeBERTa pretrained and finetuned on MNLI, SOLAR-10.7B quantized from TheBloke (for synthetic data generation)
-Encoder and classifier, fine-tuned in various ways (including with synthetic data)
+SHROOM model-agnostic dataset, DeBERTa pretrained and finetuned on MNLI, SOLAR-10.7B quantized from TheBloke (for synthetic data generation) Encoder and classifier, fine-tuned in various ways (including with synthetic data)
 MARiA
 Sanayei et al. (2024)
-SHROOM dataset, SBERT, bart-large-mnli, Mixtral
-Three approaches: (1) Cosine similarity of SBERT embeddings between source-hypothesis and source-target pairs; (2) NLI classification using bart-large-mnli model; and (3) Mixtral prompting.
+SHROOM dataset, SBERT, bart-large-mnli, Mixtral Three approaches: (1) Cosine similarity of SBERT embeddings between source-hypothesis and source-target pairs; (2) NLI classification using bart-large-mnli model; and (3) Mixtral prompting.
 Only the Mixtral results were submitted.
@@
 Bahad et al. (2024)
-SHROOM dataset; Mixtral and RoBERTa models.
-Mixtral prompting and RoBERTa finetuning.
+SHROOM dataset; Mixtral and RoBERTa models. Mixtral prompting and RoBERTa finetuning.
 NU-RU
 Markchom et al. (2024)
-SHROOM, GPT-3.5, Sentence Transformers
-Tried two approaches: (1) hypothesis-target cosine similarity, using a threshold value to determine whether the hypothesis is a hallucation.
+SHROOM, GPT-3.5, Sentence Transformers Tried two approaches: (1) hypothesis-target cosine similarity, using a threshold value to determine whether the hypothesis is a hallucation.
 (2) SelfCheckGPT with a customized prompt for each NLG task, designed to assess its coherence with the provided source and target.
@@
 Brodoceanu (2024)
-RoBERTa
-Used a pretrained model (roberta-large-openai-detector) that has been trained to distinguish between text generated by LLMs and text written by humans.
+RoBERTa Used a pretrained model (roberta-large-openai-detector) that has been trained to distinguish between text generated by LLMs and text written by humans.
 OPDAI
 Chen et al. (2024)
-SHROOM, Mistral-7B-Instruct-v0.2, self constructed training data
-Supervised fine-tuning over synthetically constructed weakly supervised training data.
+SHROOM, Mistral-7B-Instruct-v0.2, self constructed training data Supervised fine-tuning over synthetically constructed weakly supervised training data.
 Pollice Verso
@@
 Allen et al. (2024)
-SHROOM dataset; GPT 3.5 and GPT 4 models.
-In-context learning with role-play and automatic prompt generation in a few-shot classfier, using a closed-source LLM.
-SibNN
-SHROOM datasets; XLM-RoBERTa model.
-Fine-tunes a self-adaptive hierarchical variant of XLM-RoBERTa-XL twice: first as an embedder (in a few-shot mode), then as a binary classifier. More details at
+SHROOM dataset; GPT 3.5 and GPT 4 models. In-context learning with role-play and automatic prompt generation in a few-shot classfier, using a closed-source LLM.
+SibNN SHROOM datasets; XLM-RoBERTa model.
+Fine-tunes a self-adaptive hierarchical variant of XLM-RoBERTa-XL twice: first as an embedder (in a few-shot mode), then as a binary classifier.
+More details at
 https://huggingface.co/bond005/xlm-roberta-xl-hallucination-detector.
-silk_road
-SHROOM datasets; Vectara model.
-Fine-tunes an off-the-shelf Cross-Encoder hallucination evaluation model.
-Skoltech
-(No report)
+silk_road SHROOM datasets; Vectara model. Fine-tunes an off-the-shelf Cross-Encoder hallucination evaluation model.
+Skoltech (No report)
 SLPL SHROOM
 Fallah et al. (2024)
-SHROOM datasets; LaBSE, DeBERTa, Zephyr, Mistral and Llama2 models.
-Using two LLMs to classify and explain their decision and another LLM to judge and decide based on those explanations.
+SHROOM datasets; LaBSE, DeBERTa, Zephyr, Mistral and Llama2 models. Using two LLMs to classify and explain their decision and another LLM to judge and decide based on those explanations.
 SmurfCat
@@
 However, there are two other systems: one works as a voting ensemble of multiple LLMs, and another uses the Mutual Implication Score architecture.
-Team CentreBack
-SHROOM dataset; DeBERTa model.
+Team CentreBack SHROOM dataset; DeBERTa model.
 Uses an off-the-shelf library (SelfCheckGPT’s SelfCheckNLI function) to calculate contradiction scores on a small labeled test set and then defined a threshold for hallucination.
@@
 The best model-agnostic track submission is a meta-model that utilizes linear regression and is trained on features that correspond to probabilities predicted by individual systems we implemented.
-UCC-NLP
-SHROOM dataset; GPT-3.5 and Vectara models.
-Uses BertScore and GPT-3.5 to create synthetic labels and fine-tune a Vectara LLM.
+UCC-NLP SHROOM dataset; GPT-3.5 and Vectara models. Uses BertScore and GPT-3.5 to create synthetic labels and fine-tune a Vectara LLM.
 UMUTeam
 Pan et al. (2024)
-SHROOM dataset; TULU-DPO model.
-Zero-shot approach
-ustc_xsong
-(No report)
-zhuming
-(No report)
+SHROOM dataset; TULU-DPO model. Zero-shot approach
+ustc_xsong (No report)
+zhuming (No report)
 0x.Yuan
@@
 Mistral, Mixtral, LLaMA, Falcon, WizardLM and Capybara models.
-Zero-shot prompt engineering. Expects most LLMs will have different hallucination patterns, and tests whether ensembling can mitigate this.
+Zero-shot prompt engineering.
+Expects most LLMs will have different hallucination patterns, and tests whether ensembling can mitigate this.
 Table 2: Participating teams and their respective works.